# CCC++: Optimized Color Classified Colorization with Segment Anything Model (SAM) Empowered Object Selective Color Harmonization

### v1

@@ IV-C3 Quantitative Comnparison
In Tab. IX, we evaluate our proposed model against those methods across three datasets using LPIPS and FID criteria. The table shows that our method performs well in all datasets and outperforms others in the ‘Oxford Flower’ dataset. Because ‘Oxford flowers’ dataset has the highest diversity compared to ‘ADE’ and ‘Celeba.’
@@
In Tab. X, we evaluate our proposed model against those methods across five dataset using CNR criteria. The table shows that our method outperforms all methods in all the datasets. The main objective of our proposed model is to ensure the presence of minor colors along with major colors. Minor color confirmation makes color images more diverse because an image contains one or two major colors as well as more minor colors.
@@
In Tab. VI, quantitatively evaluate the images of Fig. 14. From the visual and quantitative analysis, we find Deoldify[50] has the best SSIM, Iizuka[24] has the best MSE and PSNR, Larsson[20] has the best UIQI, CIC[30] has the best LPIPS, DD[65] has the best FID, and CCC++ has the best CNR, CCAR, and TAR. Visually, the CCC++ has a more plausible major and minor color combination than the others. Therefore, it is evident that MSE, PSNR, SSIM, LPIPS, and FID criteria are not completely suitable for ensuring the presence of minor colors. Initially prioritizing the optimization of CNR, CCAR, and TAR, followed by the consistent adherence to conventional criteria, emerges as the most effective approach to enhance the representation of both major and minor colors in the generated images.
@@ IV-D Ancient Black and White Photos Colorization
Gray
CCC++
Migrant Mother,1936 Winston Churchill,1941 Kent State Shootings,1970 Birmingham,Alabama,1963 Country Doctor,1948
Figure 19: Black and white historical images and their colorized version by our proposed CCC++ method.
In Tab. VII, we evaluate our proposed model against seven baselines and SOTA methods across three dataset using regression criteria. The table shows that our method performs well in all datasets and outperforms others in the ‘Oxford Flower’ dataset. Since ‘ADE’ predominantly features natural images, while ‘Celeba’ focuses on human faces, typically presenting a more limited range of color combinations. In contrast, the ‘Oxford Flowers’ dataset is characterized by its diverse array of flower species, each exhibiting a unique and varied color palette. This diversity in coloration within the ‘Oxford Flowers’ dataset provides a more complex and challenging environment for colorization, highlighting the efficacy of our method in handling a wide range of colors and complexities.
In Tab. VIII, we evaluate our proposed model against those methods across three dataset using similarity measurement criteria. The table shows that our method performs well in all datasets while maintaining the minor color structure. Usually, it is easier to achieve good similarity by ignoring the minor color features and focusing only on major ones. However, our proposed method maintains satisfactory similarity while ensuring the minor color features.
In the presentation of five historical images alongside their meticulously colorized counterparts generated by our model in Figure 19, a compelling narrative of visual eloquence emerges. These archival photos colors. Our model’s colorization proficiency extends beyond mere aesthetics, demonstrating a keen understanding of historical contexts and nuances. The sepia-toned warmth of vintage photographs and the intricate interplay of hues in landscapes from bygone eras attest to the model’s ability to authentically recreate the colors of yesteryear. This adept infusion of diverse colors not only revitalizes the historical imagery but also serves as a poignant bridge across time, offering a contemporary perspective on moments captured in the past. The seamless integration of varied hues stands as a testament to the nuanced capabilities of our model, fostering a deeper and more resonant connection to the historical narrative embedded within each image.
@@ IV-E Ablation Study
In this section, we rigorously conduct an ablation analysis to systematically dissect and assess the distinct contributions of pivotal components within our proposed model. Illustrated in Figure 13, we offer a comprehensive comparison across various bin sizes, examining class points, trainable parameters, and average deviation. The figure reveals an intriguing trend wherein an increase in bin size corresponds to a decrease in class points and a concurrent rise in average deviation. While bin size 4 exhibits exceptionally high parameters, it also incurs elevated class points. Conversely, other bin sizes demonstrate a trade-off between parameters, class points, and average deviation. After meticulous evaluation, we strategically choose bin size 6 for our proposed model. Despite lower deviations observed in bin size 4, the associated surge in class points and parameters renders it suboptimal. Conversely, although parameters and class points decrease in other bin sizes, a simultaneous increase in average deviation diminishes their viability. In Figure 15, we present a visual representation featuring six distinct images across all bin sizes alongside their grayscale and ground truth counterparts. This visualization underscores the noteworthy correlation between lower bin sizes and heightened saturation, affirming the efficacy of our bin size selection in yielding plausible and vibrant colorization across diverse images.
In Figure 17, we present the results for six different images across seven distinct minimum appearance value determining hyperparameter ($\mathcal{P}$) selection percentages, accompanied by their grayscale and ground truth counterparts. A discernible trend emerges, showcasing that lower percentage selections result in heightened saturation. Specifically, a minimum bin selection of 4% produces a more aggressive and vivid color palette, while the 16% selection yields a softer and less saturated colorization. Since low percentage make makes the minimum appearance value little, the minor start domination over minor. Likewise more percentage makes the minimum appearance of minor value large and the weights of the minor classes increase very little. Striking a balance between saturation and subtlety, we strategically choose a minimum bin selection percentage of 10%. This choice is substantiated by the visual evidence in the figure, where the 10% minimum bin selection consistently delivers colorization characterized by a harmonious blend of saturation, softness, and plausibility, bringing the colorized images closer to the ground truth representations.
In Figure 18, we present a visual comparison featuring five images both with and without color harmonization, alongside their grayscale and ground truth counterparts. It is crucial to note a nuanced aspect here: during the introduction of minor colors into the images, occasional inconsistencies may arise. While such inconsistencies are not ubiquitous, their occasional occurrence necessitates a solution. The figure illustrates instances of color bleeding and inconsistencies in the model output, which are effectively addressed through the application of our color harmonization method. The visual transformation post-harmonization renders the images exceptionally plausible, eliminating inconsistencies and enhancing the overall aesthetic appeal. This nuanced improvement underscores the efficacy of our color harmonization technique in refining the output and ensuring a more coherent and visually pleasing result.
@@ V Conclusion
Automatically colorizing grayscale photographs with objects of diverse colors and sizes presents a formidable challenge due to the intricate variations both within and between objects, compounded by the limited spatial coverage of principal elements. The learning process often grapples with favoring dominant features, thereby yielding biased models. To mitigate this, we introduce a weighted function designed to address feature imbalance, affording greater importance to minority features. In this paper, we put forth a set of formulas facilitating the conversion of color values into corresponding color classes and vice versa. We experiment on different bin size and propose 6 as optimal bin size value. To achieve ideal performance, we strategically optimize the class levels to 532 by experimenting on a large scale of diverse category images and establish a judicious trade-off between the weights assigned to major and minor classes, ensuring accurate class prediction for both. We propose a hyperparameter in the weightening formula to make a balance relation between the weights of the major and minor classes. Additionally, we introduce SAM-empowered object-selective color harmonization, a novel approach enhancing the stability of minor classes. Our contribution extends to proposing two innovative color picture assessment measures, namely Color Class Activation Ratio (CCAR) and True Activation Ratio (TAR), complementing our previously introduced Chromatic Number Ratio (CNR), to quantitatively evaluate color component richness. Extensive evaluations against eight baseline and state-of-the-art models, conducted on six datasets, Place, ADE, Celeba, COCO, Oxford 102 Flower, and ImageNet, reveal that our proposed model outperforms its predecessors across various criteria, including visualization, CNR, CCAR, and TAR, while maintaining commendable performance in regression criteria such as MSE, PSNR, similarity criteria SSIM, LPIPS, UIUI, and generative criteria FID.

### v2

@@ IV-C3 Quantitative Comnparison
Quantitatively evaluate the images shown in Fig. 14 by referring to TABLE VI. Through both visual and quantitative investigation, we have determined that Deoldify[60] achieves the highest Structural Similarity Index (SSIM) compared to other methods, such as Iizuka[28] has achieved the highest Mean Squared Error (MSE) and Peak Signal-to-Noise Ratio (PSNR) among all the competitors. This information is supported by the research conducted by Larsson et al. cite Larsson achieves the highest UIQI score, as reported by Zhang et al. in their ECCV paper, CIC[33] achieves the highest LPIPS score, as reported by Zhang et al. in the same paper. DD achieves the highest FID score, as reported in the paper referenced as [41]. CCC++ achieves the highest CNR, CCAR, and TAR scores. The CCC++ exhibits a visually more convincing blend of major and minor colors compared to the others. Hence, it is apparent that MSE, PSNR, SSIM, LPIPS, and FID metrics are not entirely appropriate for guaranteeing the detection of subtle hues. The most effective strategy to enhance the depiction of both main and minor colors in the generated images is to initially prioritize the optimization of CNR, CCAR, and TAR, followed by consistent adherence to conventional standards. In TABLE VII, we assess the performance of our proposed model compared to seven baseline models and state-of-the-art approaches on three datasets using regression criteria. The data presented in the table demonstrates the strong performance of our method across all datasets, particularly in the ’Oxford Flower’ dataset, where it outperforms other methods. Given that ‘ADE’ mostly demonstrates natural photos and ‘Celeba’ specifically emphasizes human faces, it is common for ‘Celeba’ to exhibit a narrower variety of color combinations than ‘ADE’. On the other hand, the ‘Oxford Flowers’ dataset stands out for its wide range of flower species, each displaying a distinct and diversified color scheme. The presence of many colors and complexity in the ‘Oxford Flowers’ dataset creates a more intricate and demanding colorization setting, showcasing our method’s effectiveness in dealing with a diverse range of colors and complexities. We assess our proposed model against other approaches on three datasets using similarity measurement criteria, as shown in TABLE VIII. The table demonstrates the strong performance of our technique across all datasets while preserving the subtle color structure. Typically, it is more straightforward to attain a strong resemblance by disregarding insignificant color characteristics and concentrating on prominent ones. Nevertheless, our proposed strategy preserves a reasonable level of resemblance while preserving small color details.
Gray
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
CCC++
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a)           (b)           (c)            (d)             (e)
Figure 19: Black and white historical images and their colorized version by our proposed CCC++ method. (a) Migrant Mother, 1936 (b) Winston Churchill, 1941 (c) Kent State Shootings, 1970 (d) Birmingham,Alabama, 1963 (e) Country Doctor, 1948.
We assess our proposed model against other approaches on three datasets using the LPIPS and FID criteria, as shown in TABLE IX. The data presented in the table demonstrates the strong performance of our method across all datasets, particularly in the ’Oxford Flower’ dataset, where it outperforms other methods. The ‘Oxford flowers’ dataset exhibits greater diversity than the ‘ADE’ and ‘Celeba’ datasets. In TABLE X, we assess the effectiveness of our proposed model compared to other techniques on five datasets using the CNR, CCAR, and TAR criteria. The data presented in the table demonstrates that our approach surpasses all other methods across all datasets. The primary goal of our suggested approach is to guarantee the inclusion of both minor and major colors. Confirming minor colors enhances the diversity of color representations by including other colors alongside one or two dominant colors.
@@ IV-D Ancient Black and White Photos Colorization
The presentation exhibits five historical photos with their colorized equivalents, created by our model, as seen in Fig. 19. This pair of images creates a captivating tale of visual eloquence. The colorization skill of our model beyond basic aesthetics, featuring a deep comprehension of historical contexts and subtleties. The model’s capacity to accurately replicate the colors of the past is evident in the nostalgic sepia tones of historical images and the delicate blending of hues in landscapes from previous eras. This skillful use of various colors not only rejuvenates the historical visuals but also acts as a meaningful connection between different time periods, providing a modern viewpoint on filmed events from the past. The flawless incorporation of diverse shades serves as evidence of the subtle capabilities of our model, promoting a more profound and impact connection to the historical story inherent in each image.
@@ IV-E Ablation Study
In this section, we rigorously conduct an ablation analysis to systematically dissect and assess the distinct contributions of pivotal components within our proposed model. Illustrated in Fig. 13, we offer a comprehensive comparison across various bin sizes, examining class points, trainable parameters, and average deviation. The figure reveals an intriguing trend wherein an increase in bin size corresponds to a decrease in class points and a concurrent rise in average deviation. While bin size 4 exhibits exceptionally high parameters, it also incurs elevated class points. Conversely, other bin sizes demonstrate a trade-off between parameters, class points, and average deviation. After meticulous evaluation, we strategically choose bin size 6 for our proposed model. Despite lower deviations observed in bin size 4, the associated surge in class points and parameters renders it sub-optimal. Conversely, although parameters and class points decrease in other bin sizes, a simultaneous increase in average deviation diminishes their viability. In Fig. 15, we present a visual representation featuring six distinct images across all bin sizes alongside their grayscale and ground truth counterparts. This visualization underscores the noteworthy correlation between lower bin sizes and heightened saturation, affirming the efficacy of our bin size selection in yielding plausible and vibrant colorization across diverse images. In Fig. 17, we present the results for six different images across seven distinct minimum appearance value determining hyperparameter ($\mathcal{P}$) selection percentages, accompanied by their grayscale and ground truth counterparts. A discernible trend emerges, showcasing that lower percentage selections result in heightened saturation. Specifically, a minimum bin selection of 4% produces a more aggressive and vivid color palette, while the 16% selection yields a softer and less saturated colorization. Since low percentage make makes the minimum appearance value little, the minor start domination over minor. Likewise more percentage makes the minimum appearance of minor value large and the weights of the minor classes increase very little. Striking a balance between saturation and subtlety, we strategically choose a minimum bin selection percentage of 10%. This choice is substantiated by the visual evidence in the figure, where the 10% minimum bin selection consistently delivers colorization characterized by a harmonious blend of saturation, softness, and plausibility, bringing the colorized images closer to the ground truth representations. In Fig. 18, we present a visual comparison featuring five images both with and without color harmonization, alongside their grayscale and ground truth counterparts. It is crucial to note a nuanced aspect here: during the introduction of minor colors into the images, occasional inconsistencies may arise. While such inconsistencies are not ubiquitous, their occasional occurrence necessitates a solution. The figure illustrates instances of color bleeding and inconsistencies in the model output, which are effectively addressed through the application of our color harmonization method. The visual transformation post-harmonization renders the images exceptionally plausible, eliminating inconsistencies and enhancing the overall aesthetic appeal. This nuanced improvement underscores the efficacy of our color harmonization technique in refining the output and ensuring a more coherent and visually pleasing result.
@@ V Conclusion
Automatically colorizing grayscale images with objects of diverse colors and sizes presents a formidable challenge due to the intricate variations both within and between objects, compounded by the limited spatial coverage of principal elements. The learning process often grapples with favoring dominant features, thereby yielding biased models. To mitigate this, we introduce a weighted function designed to address feature imbalance, affording greater importance to minority features. In this paper, we put forth a set of formulas facilitating the conversion of color values into corresponding color classes and vice versa. We experiment on different bin size and propose 6 as optimal bin size value. To achieve ideal performance, we strategically optimize the class levels to 532 by experimenting on a large scale of diverse category images and establish a judicious trade-off between the weights assigned to major and minor classes, ensuring accurate class prediction for both. We propose a hyperparameter in the weightening formula to make a balance relation between the weights of the major and minor classes. Additionally, we introduce SAM-empowered object-selective color harmonization, a novel approach enhancing the stability of minor classes. Our contribution extends to proposing two innovative color picture assessment measures, namely Color Class Activation Ratio (CCAR) and True Activation Ratio (TAR), complementing our previously introduced Chromatic Number Ratio (CNR), to quantitatively evaluate color component richness. Extensive evaluations on six datasets—Places, ADE, CelebA, COCO, Oxford 102 Flower, and ImageNet—against eight baseline and state-of-the-art models demonstrate that our proposed model consistently outperforms its predecessors. It excels across multiple criteria, including visualization, CNR, CCAR, and TAR while maintaining strong performance in regression metrics such as MSE and PSNR, similarity like including SSIM and LPIPS, as well as generative metrics like the UIQI and FID.